OpenStudy (anonymous):

please helpppp Violet obtained a 25-year, fixed-rate mortgage for $176,900 on a home that cost her $194,300. If the interest rate on the mortgage is 4.82%, how much are Violet’s monthly payments? $1,015.67 $1,115.58 $8,526.58 $9,365.27

OpenStudy (anonymous):

@ankit042

OpenStudy (amistre64):

i havent figured a good way to deal with the IRA stuff ....

OpenStudy (anonymous):

Ok can you help with another then?

OpenStudy (amistre64):

P = Bk^(12*25) (1-k)/(1-k^(12*25)), k=1+.0482/12 B = 176900

OpenStudy (anonymous):

wait so what is the answer? @amistre64

OpenStudy (anonymous):

omg thank you so much

OpenStudy (amistre64):

good luck